OpenStudy (anonymous):

A loon flying horizontally at an altitude of 3 meters and speed of 8 m/s passes over a turtle standing on the ground. Find the rate at which the distance between them is changing when they are 5 meters apart

OpenStudy (anonymous):

Use \[ z^2 = x^2 + y^2 \]In this case \(y=3\) and \(y'=0\)

OpenStudy (anonymous):

Implicit differentiation gives you: \[ 2zz' = 2xx' \]We also know that \(x' = 8\)

OpenStudy (anonymous):

Also \(z=5\) and they want you to find \(z'\)

OpenStudy (anonymous):

\(x\) by Pythagorean theorem is just: \(\sqrt{5^2-3^2}\)

OpenStudy (anonymous):

@Mer94 Get it?